### v3

# ChineseHarm-Bench: A Chinese Harmful Content Detection Benchmark

Abstract
Large language models (LLMs) have been increasingly applied to automated harmful content detection tasks, assisting moderators in identifying policy violations and improving the overall efficiency and accuracy of content review. However, existing resources for harmful content detection are predominantly focused on English, with Chinese datasets remaining scarce and often limited in scope. We present a comprehensive, professionally annotated benchmark for Chinese content harm detection, which covers six representative categories and comprises a total of 6,000 samples drawn entirely from real-world data. Our annotation process further yields a knowledge rule base that provides explicit expert knowledge to assist LLMs in Chinese harmful content detection. In addition, we propose a knowledge-augmented baseline that integrates both human-annotated knowledge rules and implicit knowledge from large language models, enabling smaller models to achieve performance comparable to state-of-the-art LLMs [^1] https://github.com/zjunlp/ChineseHarm-bench..
Introduction
Harmful content detection plays a critical role in maintaining a civilized social media platform  (Jiawen et al. 2022; Jahan and Oussalah 2023; Xiao, Bouamor, and Zaghouani 2024a). The unchecked circulation of harmful or illicit content can lead to severe societal, psychological, and legal consequences (Thomas et al. 2021; Guo et al. 2024). With the massive scale of online data rendering manual detection infeasible, recent research has increasingly focused on leveraging LLMs for automated harmful content detection (Thomas et al. 2021; Guo et al. 2024; He et al. 2024; Zhang et al. 2024a; Kang and Qian 2024). Nevertheless, the majority of existing benchmarks and datasets for harmful content detection are focused on English, with Chinese resources remaining scarce and limited in scope (Xu et al. 2020; Wang et al. 2024; Yang et al. 2025). Even when Chinese datasets are available, they typically focus on a single violation category, most commonly hate speech, and thus fail to capture the full spectrum of content safety challenges encountered on Chinese platforms (Jiawen et al. 2022; Lu et al. 2023; Xiao et al. 2024; Bai et al. 2025; Yang et al. 2025).
Harmful content detection presents unique challenges that extend beyond those addressed by traditional NLP tasks (Tobi 2024). In particular, the Chinese language is highly complex and exhibits unique linguistic characteristics (Xu et al. 2023; Fang et al. 2025), further complicating harmful content detection in Chinese online environments. There exist a wide variety of perturbation methods in Chinese for evading detection, such as the use of homophones, homographs, and other similar strategies (Su et al. 2022; Xiao et al. 2024). For example, as illustrated in Figure 1 under the Abuse category, users may replace the keyword mother with a homophonic word such as piano, whose Chinese transliteration shares a similar pronunciation, in order to circumvent detection.
Figure 1: The six categories of our ChineseHarm-Bench and corresponding example cases.
To address these gaps, we present ChineseHarm-Bench, a comprehensive multi-category benchmark designed for Chinese harmful content detection. ChineseHarm-Bench is constructed from real-world violation records and covers six representative categories: gambling, pornography, abuse, fraud, illicit advertisements, and non-violation. Notably, every text and label in our benchmark has been validated by professional annotators, guaranteeing high quality and reliability. Moreover, our annotation process yields a knowledge rule base that can be used as an external knowledge source to guide human annotators and support LLMs in automated harmful content detection.
LLMs rely on pretraining data, which remains static once training is complete, limiting their ability to adapt to new or evolving information (Bigoulaeva, Madabushi, and Gurevych 2025). Since pretraining data is typically clean and curated, it may lack comprehensive coverage of harmful content, which evolves dynamically and often exhibits adversarial patterns. To address these limitations and improve resource efficiency (Bai et al. 2024; Wang et al. 2025), we introduce a knowledge-augmented baseline (Zhu et al. 2025) that enhances the performance of smaller LLMs for Chinese harmful content detection. Incorporating external knowledge, such as human-annotated rule bases, provides up-to-date priors that help both annotators and models recognize subtle violations. By constructing diverse synthetic detection scenarios through structured prompt design (Markov et al. 2022; Yu et al. 2023; Chen et al. 2022) and leveraging both explicit rules and teacher-generated responses during training, our approach enables smaller LLMs to achieve performance comparable to state-of-the-art large models while maintaining efficiency and accessibility.
Our main contributions are as follows:
  • We present a multi-category, professionally annotated benchmark for Chinese harmful content detection, which can be used to evaluate the detection capabilities of LLMs in handling harmful content in Chinese contexts.
  • We manually construct a content safety knowledge rule base during the annotation process, which not only facilitates future annotation efforts but also serves as external knowledge to enhance model detection capabilities.
  • We propose a knowledge-augmented baseline, and extensive experiments demonstrate that incorporating external knowledge allows relatively small models to achieve detection performance on par with state-of-the-art LLMs.
Related Works
Content Harm Detection.
Automated content safety detection plays a crucial role in enhancing community security (Schmidt and Wiegand 2017; Xiao, Bouamor, and Zaghouani 2024b). Initially, methods such as keyword-based detection and topic analysis were employed to identify unsafe content (Warner and Hirschberg 2012; MacAvaney et al. 2019; Deng et al. 2022). Subsequently, smaller models such as BERT (Devlin et al. 2019a) have been employed and trained on various datasets for the task of harmful content detection (Wulczyn, Thain, and Dixon 2016; Zampieri et al. 2019; Jiawen et al. 2022; Markov et al. 2022). Owing to the exceptional capabilities of LLMs, there has been a rise in approaches that directly utilize these models for harmful content detection (Guo et al. 2023; He et al. 2023; Huang, Kwak, and An 2023; Zhang et al. 2024b). Moreover, a series of guard models have recently emerged, specifically designed for harmful content detection (Inan et al. 2023; Team 2024a; Llama Team 2024; Ma et al. 2024; Zeng et al. 2024; Zhang et al. 2024d; Wen et al. 2025). However, these models primarily focus on English content and are concerned with the output safety of large models, which differs from the content safety definitions specific to the Chinese internet.
Chinese Resources.
Over the years, several datasets have been proposed to address specific aspects of harmful content detection in the Chinese language. COLA (Tang and Shen 2020) provides the first Chinese offensive language classification dataset, while SWSR (Jiang et al. 2022) introduces the first Chinese dataset specifically targeting sexist content. COLD (Jiawen et al. 2022) provides a nuanced taxonomy of Chinese offensive content, while TOXICN (Lu et al. 2023) expands toxicity detection to both explicit and implicit cases. Building on this, ToxiCloakCN (Xiao et al. 2024) introduces perturbed examples to evaluate model robustness. Furthermore,  Bai et al. (2025) present a span-level toxicity extraction dataset, and SCCD (Yang et al. 2025) offers fine-grained comment-level annotations for Chinese cyberbullying detection. Despite these advancements, the focus of these datasets predominantly remains on hate speech, whereas the scope of Chinese content detection extends beyond this singular aspect. Recent datasets such as SafetyBench (Zhang et al. 2024c) and ChineseSafe (Zhang et al. 2025) have attempted to address broader categories of harmful content. However, some categories in these datasets, while related to other aspects of safety, are not directly relevant to harmful content detection, as certain types of content are considered acceptable and can be freely circulated on Chinese platforms.
Benchmark
Figure 2: Overview of the benchmark construction pipeline. The process includes data collection and filtering, clustering-based sampling, and expert annotation with iterative knowledge rule base refinement. Finally, 1,000 instances are sampled for each category to form the final benchmark.
Figure 2 illustrates the overall construction process of our benchmark. We collect and filter real-world data, perform clustering-based sampling, and conduct expert annotation with iterative knowledge rule base refinement. This pipeline ensures a balanced, high-quality dataset with explicit knowledge rules for each category.
Benchmark Category
Based on Chinese laws and regulations [^2] https://www.gov.cn/gongbao/content/2000/content˙60531.htm [^3] https://www.gov.cn/gongbao/content/2020/content˙5492511.htm, we select six representative categories for our study: Gambling, Pornography, Abuse, Fraud, Illicit Ads, and Non-violation. These categories cover a broad range of application scenarios and demonstrate strong representativeness and research value. The detailed selection rationale is provided in Appendix A.
Figure 1 provides a simple example and its translation for each category. The basic definitions of these six categories are as follows:
  • Gambling: Content related to gambling activities, including promotion of betting platforms, sharing gambling experiences, or encouraging participation. Gambling is strictly prohibited by Chinese law due to risks of financial loss, addiction, family disruption, and social instability.
  • Pornography: Content containing vulgar or obscene material related to sexual acts, such as explicit descriptions, images, or videos. Dissemination of pornographic content is illegal in China, as it undermines social morals, harms minors, and disrupts public order.
  • Abuse: Content involving abusive language, insults, or provocation, including personal attacks, hate speech, or harassment. Such content is prohibited by Chinese regulations as it can cause psychological harm, disrupt social harmony, and incite violence or discrimination.
  • Fraud: Content involving deceptive practices intended to mislead or defraud, such as phishing, scam advertisements, or impersonation. Fraud is a criminal offense under Chinese law, posing risks to property and information security, and undermining social trust.
  • Illicit Ads: Content advertising illegal activities or products, including unlicensed drugs, counterfeit goods, or prohibited services. Publishing illicit advertisements is strictly forbidden, as it facilitates criminal activity, endangers public safety, and violates consumer rights.
  • Non-violation: Content that complies with Chinese laws and regulations and does not fall into the above categories. Such content is considered legal and appropriate for dissemination in China.
Data Collection
Data Source.
Our violation data is sourced from one of the largest social platforms in China. We collected real-world violation records from the internal database of an online platform over recent years, covering the five categories described above [^4] Due to anonymity requirements, we do not disclose the platform’s name.. Each data instance is represented as a tuple $x=(\text{text},\text{label})$, where text denotes the message content and $\text{label}\in\mathcal{C}$ is the corresponding category label. The original records are annotated with a single label upon collection, and each violation category contains approximately 15,00015{,}000 samples. Non-violation data is sourced from the Alpaca-Chinese (Taori et al. 2023; Ziang Leng and Li 2023) dataset, which provides approximately 52,00052{,}000 diverse and legally compliant responses.
Preliminary Processing.
Due to the proprietary nature of the platform’s internal annotation guidelines, some labels may be inaccurate and not all annotations have undergone thorough manual review. Given the impracticality of fully manual annotation, we designed a data filtering and optimization pipeline to ensure data quality and diversity. Specifically, we first deduplicate the data within each category. Then, for each category $c\in\mathcal{C}$, we apply k-means clustering with $100$ clusters on sentence embeddings generated by bert-base-chinese. From each cluster, we randomly sampled $20$ instances, resulting in a benchmark set of 2,0002{,}000 samples per category.
Human Annotation
Annotator Background.
To ensure annotation quality, we recruited three professional annotators from a specialized annotation team. All annotators are native Chinese speakers (two males and one female), with two holding bachelor’s degrees and one holding an associate degree. Each annotator has substantial prior experience in data annotation and harmful content detection, ensuring familiarity with relevant legal and ethical considerations. Before the annotation process, all annotators received additional training on the specific task requirements and labeling criteria to further standardize the annotation process.
Annotation Assignment and Labeling Protocol.
Each annotator was assigned responsibility for two specific categories to ensure focused expertise and consistency within categories. Since the candidate samples already carried preliminary labels from prior internal processes, the annotation mainly functioned as a verification step, which considerably reduced the annotation complexity. Given this relatively straightforward task, and considering practical constraints such as limited budget and time, we did not implement multi-annotator labeling or redundancy. Instead, we relied on experienced annotators and iterative refinement of standardized guidelines to uphold annotation quality.
Annotator Training and Calibration.
Annotators first participated in comprehensive training sessions covering the annotation guidelines, task objectives, and detailed category definitions. Annotators were also instructed on how to iteratively refine the knowledge rule base throughout the annotation process. Prior to the formal annotation, a multi-round calibration phase was conducted. Each annotator labeled 100 randomly selected samples per category, followed by alignment discussions with the authors to resolve ambiguities and ensure consistency. This procedure was repeated for three rounds until sufficient agreement was achieved, after which annotators proceeded to the full annotation stage.
Annotation Process.
Let $\mathcal{D}_{c}=\{x_{i,c}\}_{i=1}^{N}$ denote the set of N=2, [^0] N=2{,}000 candidate samples for category $c$. We initialized the knowledge rule base $\mathcal{R}_{c}=\emptyset$ for each category. The annotation process proceeds as follows for each sample $x_{i,c}$:
  • If $x_{i,c}$ matches any rule $r\in\mathcal{R}_{c}$, we retain $x_{i,c}$ in $\mathcal{D}_{c}$.
  • If $x_{i,c}$ truly belongs to category $c$ and does not match any rule in $\mathcal{R}_{c}$, we update an existing rule or add a new rule $r_{i,c}$ to $\mathcal{R}_{c}$, and retain $x_{i,c}$ in $\mathcal{D}_{c}$.
  • If $x_{i,c}$ does not belong to category $c$, we discard $x_{i,c}$ from $\mathcal{D}_{c}$.
After this process, we randomly sampled M=1, [^0] M=1{,}000 instances from the retained set for each category to ensure class balance. This procedure guarantees that our final dataset is both diverse and balanced, with all samples annotated by human experts. Furthermore, we iteratively refined the standardized annotation guideline knowledge rule base $\mathcal{R}=\bigcup_{c\in\mathcal{C}}\mathcal{R}_{c}$ (see Appendix Table 6).
Evaluation Metrics
Our benchmark is designed to evaluate the Chinese harmful content detection capabilities of LLMs. Specifically, we adopt a zero-shot setting, where the model is prompted to classify each input instance into one of the predefined categories using a standardized template (see Appendix C). Given a content item to be detected, we construct the model input as:
$X=\mathrm{Prompt\_Detect}(\mathcal{R},\,\mathrm{content})$ (1)
where $\mathcal{R}$ denotes the human-annotated knowledge rule and $\mathrm{content}$ represents the content item to be detected. Here, $\mathrm{Prompt\_Detect}(\cdot)$ refers to the process of formatting the input according to the prompt template, incorporating both the rule base and the content item. The model subsequently predicts the category for each input instance. For evaluation, we report both the per-category F1 scores and the macro-F1 score. As the dataset is balanced across categories, the macro-F1 is equivalent to the weighted-F1.
A Knowledge-Augmented Baseline
Hybrid Knowledgeable Prompting
To comprehensively simulate real-world harmful content detection scenarios, we first define a set of hierarchical, fine-grained attributes that characterize different types of illicit content. We formalize the prompt construction process as a mapping from a structured user-content space to a prompt space. Specifically, we decompose the scenario space into four primary components: persona features, text features, evasion tactics, and human-annotated knowledge rules. Notably, our attribute definitions incorporate both evasion tactics and external knowledge, with the aim of more closely modeling the complexities observed in real-world illicit content. Each component is further specified by a set of secondary, fine-grained attributes. For each violation category $c$, we define the structured input as
$U_{c}=\{U_{\text{persona}},U_{\text{text}},U_{\text{evasion}},U_{\text{knowledge},c}\}$ (2)
where:
  • $U_{\text{persona}}$: Information about the author, such as gender, age, occupation, education, reflecting diverse writing styles.
  • $U_{\text{text}}$: Intrinsic properties of the text, including text length, narrative perspective, and publishing platform.
  • $U_{\text{evasion}}$: Evasion strategies commonly observed in real-world scenarios, such as the use of emojis, homophones, and other techniques to circumvent detection. See Appendix B for more detailed explanations.
  • $U_{\text{knowledge},c}$: The reference to the human-annotated knowledge rule base $\mathcal{R}_{c}$ for category $c$, specifying the particular guidelines violated by the text.
Synthetic Data Curation
We construct synthetic data by first designing a comprehensive prompt template, denoted as prompt_generate (see Appendix Table 8), which encodes the diverse attributes described above. For each category $c\in\mathcal{C}$, we define each instance by its attribute set $U_{i,c}\in U_{c}$. For each $U_{i,c}$, the input prompt is constructed as
$Q_{i,c}=\mathrm{Prompt\_Generate}(U_{i,c})$ (3)
Here, $\mathrm{Prompt\_Generate}(\cdot)$ refers to the function that formats the sampled attribute set $U_{i,c}$, together with the corresponding rule base, into a prompt suitable for input to the teacher model. A teacher model $M_{T}$ is then used to generate a candidate response for each prompt $Q_{i,c}$:
$A_{i,c}=M_{T}(Q_{i,c})$ (4)
For each category $c\in\mathcal{C}$, we collect a set of $(Q_{i,c},A_{i,c})$ pairs. To ensure data quality and class balance, we remove duplicate instances and filter out model refusals or generic non-answers using a keyword-matching strategy (see Appendix Table 7). Finally, we uniformly sample $n$ instances from each category to construct the final dataset:
$D_{\text{final}}=\bigcup_{c\in\mathcal{C}}\text{Sample}_{n}(Q_{i,c},A_{i,c})$ (5)
This pipeline ensures that the resulting synthetic dataset is diverse, high-quality, and balanced across all categories. An overview of the synthetic data curation pipeline is presented in Figure 3.
Figure 3: Overview of the synthetic data curation pipeline. We first define a set of hierarchical, fine-grained attributes to comprehensively characterize illicit content. For each category, structured prompts are constructed based on sampled user and content attributes, evasion tactics, and human-annotated knowledge rules.
Knowledge-Guided Training
To fully leverage both explicit human knowledge and implicit model knowledge, we adopt a supervised fine-tuning (SFT) framework that incorporates two distinct sources of knowledge for each training instance. Specifically, for each sample in the curated dataset $D_{\text{final}}$, we construct the input by combining (1) human-annotated knowledge $\mathcal{R}$ and (2) teacher model knowledge, represented by the answer $A_{i,c}$ generated by the teacher model, which reflects its implicit knowledge. Formally, for each entry $(Q_{i,c},A_{i,c})$ in $D_{\text{final}}$, the input to the student model is constructed as:
$X_{i,c}=\mathrm{Prompt\_Detect}(\mathcal{R},A_{i,c})$ (6)
The student model $M_{S}$ is trained to generate the target output sequence $c$. For each instance, the sequence-level loss is defined as:
$\mathcal{L}(c\mid X_{i,c},\phi)=-\sum_{t=1}^{T_{c}}\log P(c_{t}\mid X_{i,c},c_{<t};\phi)$ (7)
where $c=(c_{1},c_{2},\ldots,c_{T_{c}})$ is the tokenized category name for category $c$, and $T_{c}$ is its length. The fine-tuning objective is to minimize the average loss over all instances:
$\phi^{*}=\arg\min_{\phi}\frac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\mathcal{L}(c\mid X_{i,c},\phi)$ (8)
where $|\mathcal{C}|$ is the number of categories, $N_{c}$ is the number of instances in category $c$, $X_{i,c}$ is the input for the $i$-th instance in category $c$, and $\phi$ denotes the parameters of the student model. This training paradigm enables the student model to integrate both explicit rule-based knowledge and implicit knowledge distilled from the teacher model, thereby enhancing its detection capability.
Backbone Strategy Knowledge F1 in each category Macro-F1
Gambling Pornography Abuse Fraud Illicit ads Non-Violation
[…]
Deepseek-R1
[…]
O3-mini
[…]
GPT-4o
[…]
GPT-4o-mini
[…]
Gemini 1.5 pro
[…]
Gemini 2.0 Flash
[…]
Claude 3.5 Sonnet
[…]
Claude 3.5 Haiku
[…]
Bert-Base-Chinese
[…]
Qwen-2.5
0.5B-Instruct
[…]
Qwen-2.5
1.5B-Instruct
[…]
Qwen-2.5
3B-Instruct
[…]
Qwen-2.5
7B-Instruct
[…]
Experiments
Experimental Setup
Model Groups.
To provide a comprehensive evaluation of Chinese harmful content detection capabilities, we consider three groups of models: (1) state-of-the-art LLMs, such as Deepseek-R1 (Guo et al. 2025), GPT series (Hurst et al. 2024; Jaech et al. 2024) , Gemini series (Team et al. 2024), and Claude series (Anthropic 2024); (2) lightweight models with fewer than 1B parameters, including Bert-Base-Chinese (Devlin et al. 2019b) and the smallest Qwen-2.5 model (Team 2024b); and (3) billion-scale LLMs with 1–10B parameters, represented by a range of Qwen-2.5 models. This selection covers a wide spectrum of model sizes and architectures for Chinese harmful content detection.
Evaluation Protocol.
For models evaluated via direct prompting, we use the original, unmodified model checkpoints. For models evaluated via fine-tuning, we refer to student models trained on synthetic data generated by the teacher model.
Knowledge Augmentation.
[…]
Training and Evaluation Details.
For our proposed baseline, we use GPT-4o as the teacher model $M_{T}$ to generate synthetic data, with a temperature of 1.0 and top-$k$ sampling ($k=1$) to encourage output diversity. We sample $n=3000$ synthetic instances per category. Qwen series student models are fine-tuned using the LLaMA Factory (Zheng et al. 2024) framework. All experiments are conducted on 8 Huawei Ascend 910B NPUs (80GB each). More experimental details are provided in Appendix D.
Main Results
Current LLMs are not yet sufficient to match human annotators.
Recent LLMs have demonstrated impressive capabilities across various domains. However, as shown in Table 1, even the best-performing models, such as Deepseek-R1 and GPT-4o, achieve average macro-F1 scores of no more than 0.8 when external knowledge is incorporated, with performance dropping further in the absence of such knowledge. Additionally, deploying these models comes at the cost of significant computational resources. Smaller models, while more computationally efficient, perform even worse when used without fine-tuning, with macro-F1 scores falling below 0.5 even when external knowledge is introduced. Importantly, harmful content detection is not a static problem but a dynamic and adversarial process, where users continuously devise novel evasion strategies to bypass detection systems. These findings indicate that the task of Chinese harmful content detection remains a significant challenge for current LLMs and is still far from achieving performance comparable to that of human annotators.
Incorporating external knowledge consistently improves model performance.
As shown in Table 1, for all models with more than 1B parameters, providing human-annotated knowledge as input during direct prompting consistently yields performance improvements. The only exception is the Qwen-2.5-0.5B model, which does not benefit from external knowledge, possibly because models of this scale lack the capacity to effectively leverage complex knowledge sources. Moreover, in the fine-tuning scenario, omitting knowledge guidance during data generation leads to a significant drop in performance across all model scales. These results demonstrate that effectively incorporating external knowledge is essential to achieve optimal performance in harmful content detection tasks.
Figure 4: Left: Macro-F1 scores of student models trained on synthetic data, comparing performance with and without evasion cases. Right: Macro-F1 scores in harmful content detection, showing the relationship with the number of synthetic samples per category (x-axis in thousands).
Our knowledge-augmented approach substantially improves the performance of lightweight and billion-scale models.
As shown in Table 1, all fine-tuned models with knowledge augmentation achieve macro-F1 scores above 0.7, compared to original scores below 0.5. Notably, the Qwen-2.5-3B and Qwen-2.5-7B models reach a macro-F1 of 0.77, surpassing all state-of-the-art LLMs under direct prompting without external knowledge. This performance is also comparable to GPT-4o (0.78) and Deepseek-R1 (0.80) when provided with external knowledge. Furthermore, even with knowledge augmentation, models such as GPT-4o-mini, Claude-3.5 Sonnet, Gemini 2.0 Flash, and O3-mini do not exceed a macro-F1 of 0.77. These results demonstrate that our approach substantially enhances the harmful content detection capabilities of lightweight and billion-scale models, enabling them to achieve performance comparable to the State-of-the-Art LLMs.
Lightweight models (<1B parameters) face inherent performance ceilings.
While knowledge augmentation improves performance across all models, the lightweight models Qwen-2.5-0.5B and BERT-Base-Chinese plateau at macro-F1 scores around 0.70. In contrast, all billion-scale LLMs closely match the performance of GPT-4o when external knowledge is incorporated. These findings highlight the intrinsic limitations of sub-1B models, which remain unable to match the effectiveness of larger models on complex Chinese harmful content detection tasks, even when provided with additional knowledge.
Analysis
Effectiveness of generating evasion cases for Chinese harmful content detection.
To assess the impact of evasion cases in synthetic data, we compare models trained on data with and without evasion examples. Specifically, we used GPT-4o to generate 3k non-evasive samples per category as the baseline, keeping all other configurations unchanged. As shown in Figure 4 (left), models trained with evasion cases achieve performance gains, underscoring the importance of incorporating Chinese-specific evasion data for detection.
3,000 synthetic samples per category are sufficient for optimal performance.
To investigate the impact of synthetic data volume, we conduct experiments with 1k, 2k, 3k, and 4k samples per category. As shown in Figure 4 (right), the performance of most models generally improves as the number of synthetic samples increases, but plateaus at 3k samples per category. This suggests that using more than 3k synthetic samples per category yields diminishing returns for harmful content detection.
Using different teacher models for data generation remains effective.
We further investigate the impact of teacher model selection by using the Deeseek-R1 model for synthetic data generation, while keeping the number of samples per category at 3k and maintaining the same training setup and configurations. As shown in Table 2, our proposed baseline continues to achieve strong performance, demonstrating robustness to the choice of teacher model and highlighting the broad applicability of our approach.
Model GPT-4o DeepSeek-R1
Bert-Base-Chinese 0.70 0.69
Qwen-2.5-0.5B-Instruct 0.70 0.65
Qwen-2.5-1.5B-Instruct 0.75 0.73
Qwen-2.5-3B-Instruct 0.77 0.76
Qwen-2.5-7B-Instruct 0.77 0.76
Table 2: Macro-F1 of student models trained on synthetic data from different teacher models.
Conclusion and Future Work
In this work, we introduce a comprehensive real-world benchmark for Chinese harmful content detection, encompassing multiple violation categories and accompanied by a professionally curated knowledge rule base. We further propose a knowledge-augmented strong baseline that integrates explicit knowledge rules and implicit knowledge from large teacher models. This approach enables small models to match or even outperform much larger models, without sacrificing efficiency or accessibility. Future extensions of this work include expanding the current taxonomy to cover a broader range of violation types, thereby better reflecting the complexity of harmful content encountered in real-world settings. Introducing multi-label annotation is another promising direction, as some instances may involve multiple overlapping violation categories. Additionally, continuously updating and maintaining the knowledge rule base could further improve model generalization and robustness. We believe that ChineseHarm-Bench offers a solid foundation for advancing research on harmful content detection and contributes toward building a healthier and more trustworthy online environment.
Appendix A Category Design and Annotation Principles
Representative and Publishable Categories.
Harmful content appears in various forms, but not all types are equally relevant to detection tasks or suitable for public research. In this study, we focus on categories that are both highly representative in real-world scenarios and appropriate for open publication. Specifically, we selected the most frequent categories from a harmful content detection database and merged certain closely related types to ensure conceptual clarity and annotation consistency.
Scope Differentiation from LLM Safety.
Our task setting differs from conventional LLM safety alignment, which primarily restricts model outputs such as instructions for making explosives or generating misinformation. In contrast, our focus is on harmful content that would typically be detected and removed shortly after being published on the internet. For example, bomb-making instructions can often be easily accessed or collected online; similarly, fake news that does not spread widely is often difficult to detect in practice. These types of content, while relevant to broader discussions of safety, fall outside the scope of our benchmark due to their limited detectability and enforcement in real-world harmful content detection systems.
Single-Label Annotation Strategy.
Although a single data instance may potentially involve multiple harmful aspects, we assign only one primary category to each instance. If an instance is highly ambiguous and reasonably fits multiple categories without a clear primary one, we discard it to ensure label reliability. This design choice is based on two main considerations. First, our goal is to build a clear and practical benchmark for harmful content detection, where assigning a single label helps avoid ambiguity and ensures annotation consistency. Second, in real-world detection systems, content is often assigned a dominant category for streamlined decision-making. We prioritize the most salient or policy-relevant label for each instance, following standardized guidelines to maintain inter-annotator agreement.
Appendix B Evasion Patterns and Case Studies
Figure 5: Representative examples of four common evasion strategies: pinyin, homophones, homographs, and emoji.
On Chinese internet, users often employ various evasion strategies to circumvent harmful content detection. To simulate these real-world behaviors, we instruct the teacher model to adopt a keyword substitution strategy, systematically replacing sensitive terms with pinyin (phonetic transcription), homophones (words with similar pronunciation), homographs (visually similar), or emojis. Representative examples of these four evasion strategies are in Figure 5.
Appendix C Prompt Template for Chinese Harmful Content Detection
作为中文互联网内容审核专家,请依据中国现行法律法规及内容安全标准,判断以下文本是否适合公开发布。 如果文本违规,请指出违规类别,类别范围限定为:“博彩”、“低俗色情”、“谩骂引战”、“欺诈”、“黑产广告”。如果文本不违规,请回答“不违规”。 请直接输出一个类别:包括“博彩”、“低俗色情”、“谩骂引战”、“欺诈”、“黑产广告””或“不违规”。
规则: {规则知识} 文本: {检测文本}
Figure 6: The Prompt_Detect template for zero-shot Chinese harmful content detection. The model is instructed to classify the input text into one of five violation categories or “non-violation” according to Chinese laws and content safety standards. The 规则知识 (“rule base”) section incorporates the human-annotated knowledge rule base $\mathcal{R}$ built during our annotation process, while the 检测文本 (“input text”) section provides the text instance to be classified.
Appendix D Additional Experimental Details
For all state-of-the-art LLMs, all models except DeepSeek-R1 are accessed via APIs, while DeepSeek-R1 is deployed locally. To ensure reproducibility, we set the temperature to 0 for all API-based models. For all Qwen series models, inference is performed using greedy decoding.
For fine-tuning the Qwen series student models, we utilize the LLaMA Factory (Zheng et al. 2024) framework with the following hyperparameters: a per-device batch size of 4, gradient accumulation steps of 2, a learning rate of $1.0\times 10^{-5}$, three epochs, cosine learning rate scheduling, a warmup ratio of 0.1, and bfloat16 precision.
For the BERT-based sequence classification baseline, we employ HuggingFace Transformers library and fine-tune the model with 6 output classes. The model is trained for 3 epochs with a learning rate of $2\times 10^{-5}$, a batch size of 32 for training and 128 for evaluation, and a weight decay of 0.01. Mixed-precision (fp16) training is enabled to accelerate computation and reduce memory usage. Other hyperparameters follow default settings of the Transformers library.
Appendix E Supplementary Tables Corresponding to the Main Text
This section includes supplementary tables that provide implementation details and extended results referenced in the main text. Specifically:
  • Tables 3, 4, 5 present more detailed results corresponding to Figure 4 and Table 2 in the main text.
  • Table 6 provides the finalized human-annotated knowledge rule base $\mathcal{R}=\bigcup_{c\in\mathcal{C}}\mathcal{R}_{c}$.
  • Table 7 lists the keyword-matching strategy used to filter out model refusals and generic outputs during synthetic data construction.
  • Table 8 shows the prompt template Prompt_Generate, which incorporates diverse content attributes for each category $c\in\mathcal{C}$.
Backbone Strategy Evasion F1 in each category Macro-F1
Gambling Pornography Abuse Fraud Illicit ads Non-Violation
Bert-Base-Chinese
Finetuning w/ 0.74 0.65 0.76 0.68 0.68 0.70 0.70
Finetuning w/o 0.71 0.66 0.75 0.68 0.67 0.69 0.69
Qwen-2.5
0.5B-Instruct
Finetuning w/ 0.75 0.64 0.75 0.62 0.70 0.74 0.70
Finetuning w/o 0.69 0.60 0.75 0.63 0.58 0.63 0.65
Qwen-2.5
1.5B-Instruct
Finetuning w/ 0.77 0.71 0.77 0.70 0.74 0.79 0.75
Finetuning w/o 0.76 0.65 0.76 0.68 0.76 0.73 0.72
Qwen-2.5
3B-Instruct
Finetuning w/ 0.81 0.72 0.79 0.72 0.74 0.85 0.77
Finetuning w/o 0.80 0.68 0.82 0.72 0.76 0.76 0.75
Qwen-2.5
7B-Instruct
Finetuning w/ 0.82 0.70 0.75 0.75 0.75 0.82 0.77
Finetuning w/o 0.76 0.70 0.75 0.71 0.72 0.72 0.73
Table 3: Detailed per-category F1 and macro-F1 scores for models trained on synthetic data with and without evasion cases, corresponding to Figure 4.
Backbone Strategy Number F1 in each category Macro-F1
Gambling Pornography Abuse Fraud Illicit ads Non-Violation
Bert-Base-Chinese
Finetuning 1k 0.73 0.65 0.75 0.59 0.63 0.71 0.68
Finetuning 2k 0.72 0.64 0.75 0.60 0.67 0.66 0.67
Finetuning 3k 0.74 0.65 0.76 0.68 0.68 0.70 0.70
Finetuning 4k 0.74 0.66 0.75 0.65 0.68 0.67 0.69
Qwen-2.5
0.5B-Instruct
Finetuning 1k 0.79 0.65 0.66 0.67 0.73 0.75 0.71
Finetuning 2k 0.75 0.63 0.69 0.66 0.73 0.74 0.70
Finetuning 3k 0.75 0.64 0.75 0.62 0.70 0.74 0.70
Finetuning 4k 0.74 0.65 0.75 0.68 0.70 0.73 0.71
Qwen-2.5
1.5B-Instruct
Finetuning 1k 0.80 0.69 0.73 0.69 0.73 0.80 0.74
Finetuning 2k 0.80 0.69 0.73 0.69 0.73 0.82 0.74
Finetuning 3k 0.77 0.71 0.77 0.70 0.74 0.79 0.75
Finetuning 4k 0.80 0.71 0.75 0.69 0.72 0.80 0.75
Qwen-2.5
3B-Instruct
Finetuning 1k 0.81 0.71 0.79 0.61 0.72 0.86 0.75
Finetuning 2k 0.80 0.69 0.73 0.72 0.78 0.83 0.76
Finetuning 3k 0.81 0.72 0.79 0.72 0.74 0.85 0.77
Finetuning 4k 0.80 0.70 0.77 0.73 0.76 0.80 0.76
Qwen-2.5
7B-Instruct
Finetuning 1k 0.79 0.73 0.78 0.62 0.71 0.83 0.74
Finetuning 2k 0.79 0.68 0.75 0.67 0.72 0.81 0.74
Finetuning 3k 0.82 0.70 0.75 0.75 0.75 0.82 0.77
Finetuning 4k 0.82 0.69 0.72 0.70 0.77 0.83 0.75
Table 4: Detailed macro-F1 and per-category F1 scores for different models and numbers of synthetic samples per category, corresponding to Figure 4.
Backbone Strategy Teacher F1 in each category Macro-F1
Gambling Pornography Abuse Fraud Illicit ads Non-Violation
Bert-Base-Chinese
Finetuning GPT-4o 0.74 0.65 0.76 0.68 0.68 0.70 0.70
Finetuning DeepSeek-R1 0.77 0.67 0.75 0.65 0.61 0.70 0.69
Qwen-2.5
0.5B-Instruct
Finetuning GPT-4o 0.75 0.64 0.75 0.62 0.70 0.74 0.70
Finetuning DeepSeek-R1 0.77 0.65 0.68 0.66 0.49 0.64 0.65
Qwen-2.5
1.5B-Instruct
Finetuning GPT-4o 0.77 0.71 0.77 0.70 0.74 0.79 0.75
Finetuning DeepSeek-R1 0.82 0.72 0.77 0.73 0.66 0.72 0.74
Qwen-2.5
3B-Instruct
Finetuning GPT-4o 0.81 0.72 0.79 0.72 0.74 0.85 0.77
Finetuning DeepSeek-R1 0.82 0.75 0.77 0.77 0.74 0.73 0.76
Qwen-2.5
7B-Instruct
Finetuning GPT-4o 0.82 0.70 0.75 0.75 0.75 0.82 0.77
Finetuning DeepSeek-R1 0.84 0.73 0.80 0.76 0.74 0.71 0.76
Table 5: Detailed per-category F1 and macro-F1 scores for models trained on synthetic data generated by different teacher models, corresponding to Table 2.
Human-annotated knowledge rule
博彩: 1. 使用赌博行业术语:包括“28”、“壹号”、“问鼎”、“时时彩”、“体彩”、“发布博彩内幕”、“真人娱乐城”、“澳门娱乐城”、“加拿大”、“接龙”、“扫雷”、“红蓝”等词汇及其变体或拆分形式。“加拿大28”是一种网络赌博玩法,“接龙”和“扫雷”是微信红包赌博游戏,“壹号”和“问鼎”是博彩网站,“红蓝”通常与百家乐等赌博游戏相关。 2. 诱导点击的博彩网址:博彩网站通常使用特定的域名后缀,如“.top”和“.vip”等。 3.与赌博相关的内容:包括赌博游戏和平台(如私盘赌博、红包赌博、棋牌赌博等)、赌博行为(如开盘、投注、公布结果等)、赌博组织和人员(如赌客、财务等),以及与赌博相关的字词(如赢钱、德州扑克、老虎机等)。
低俗色情: 1. 低俗性暗示和不伦关系:使用直接或间接指向性行为、性饥渴、局部敏感部位或暗示有色情资源的关键词和语句,如“馋”、“难忍”、“需求大”、“放开”、“放纵”、“有水”、“带小雨伞”、“想吃香蕉”、“你懂得”等。包括宣扬乱伦关系如兄妹、爷孙、父女等,或描述相关性行为、性暗示、擦边行为色情裸露等内容。 2. 敏感部位和性行为的描述:使用指代词、谐音、低俗别称等对人体敏感部位进行形容和描述,对性行为、性描写、性方式、性场景进行直观展示或描述,包括对人体性部位的直接暴露和描写,涉及特殊性癖如“SM”、“绿帽癖”等。 3. 对女性外貌和气质的低俗描述:使用带有性暗示或低俗内容的语言,可能涉及贬低女性的词汇,如“母狗”、“婊子”等。 4. 色情服务和物品交易:使用隐晦或黑话传播招嫖信息,或直接发布提供色情服务的内容,包括描述身体特征的词汇如“大胸”、“爆乳”、“36D”等,或提供“上门服务”、“陪洗澡”、“按摩”等服务。涉及色情物品的买卖行为,如原味或二手内衣、袜子、真人情趣玩具等,以及传播和寻求色情资源的行为,如分享色情资源的获取手段、app、网站等,例如“91”、“吃瓜群”、“黄色网站”、“app”、“网盘链接”等。 5. 以性行为为目的的交友行为:以发生性行为、性关系为目的的交友行为,常见情形如约炮、一夜情、床伴等。
谩骂引战: 1. 人身攻击与辱骂:通过伦理身份、人格等进行攻击侮辱,编造网络黑话、恶意造梗,通过拼音、谐音、指代词等方式,恶意编造低俗烂梗、使用污言秽语侮辱谩骂他人。 2. 发布对立和歧视内容:包括性别对立、阶层对立、地域歧视等,污名化特定群体,煽动职业、性别、阶级、地域、宗族等歧视与对立,激化社会矛盾。
欺诈: 1. 高佣金兼职诈骗:以高佣金、高薪,夸大行为等诱导话术发布发布网赚、兼职任务。 2. 金融诈骗:包括股票投资诈骗,通过推荐个股或投资产品诱导他人投资;贷款诈骗,仿冒正规机构进行贷款服务欺诈;回款清退诈骗,以投资平台“清退兑付”的名义进行引流,后续进行投资诈骗。 3. 身份仿冒诈骗:仿冒公检法等国家权力机关,如刑侦、检察院、法院、银行年检专员身份诈骗、电商、快递客服等,编造理由进行欺诈。 4. 免费赠品诈骗:以活动中奖、粉丝回馈等理由免费赠送高价值礼品,包括免费领游戏皮肤,索要微信号密码、下单物品免费但需交邮费等。
黑产广告: 1. 低佣金招募闲散人员提供简单任务:包括使用“日结”、“一单一结”、“立返佣”等字眼,提供超简单任务如点赞、关注、评论、打字聊天、抄小说等,并承诺低额佣金。提供淘宝等平台的刷单、销量、知名度、增加人气等服务,或引导用户去电商平台刷单。 2. 提供账号相关服务:包括账号交易、账号代实名、代批量注册等服务,发布可通过技术手段解封社交账号、游戏账号的有偿服务内容,以及代注册、代认证服务(如帮助做人脸实名认证类)。 3. 非法金融和技术服务:提供涉及技术手段的非法操作和规避服务,包括恶意举报、一键群发、钻漏洞低价、套现、洗钱等。宣传不看信用、不看征信的借贷服务,提供非正常提额、刷信用、代开、强开服务或教程。还包括黑户漂白、信用卡逾期处理、征信修复、养卡服务,以及交易需实名的卡证信息如银行卡、信用卡、SIM卡、营业执照、对公账户等。 4. 提供接码和引流服务:包括提供接码任务(利用普通人的手机号接收验证码)或相关技术服务,利用工具人的设备、账号达到引流目的。出现收售特定群体粉丝如色情粉、赌博粉、兼职粉、交友粉等,或为黑灰产业提供引流账号、设备或技术支持等。
Table 6: Human-annotated knowledge rules for each violation category in our Chinese harmful content detection benchmark.
Filter Keywords
"I’m sorry", "Sorry", "I am sorry", "I apologize", "As an", "As an AI", "I’m an", "I’m just", "As a language model", "As an Assistant", "I cannot", "I do not", "Hello!", "is illegal and unethical", "I’m really sorry", "OpenAI", "unethical", "not ethical", "illegal", "not legal", "My apologies", "but I cannot", "If you have any other non-malicious requests", "I’m not able to provide", "I am really sorry", "I cannot create", "I cannot support", "I can’t provide", "I can’t assist", "I can’t", "I am not able", "I am unable", "I will not", "I do not engage", "illegal", "unethical", "harmful", "I am not able to"
"抱歉", "对不起", "违规内容","社区规定","无法生成","违法内容","不好意思","无法完成","违反规定","不道德","合法","不能生成","无法提供","平台规则","不符合","合规","违反法律","违规文本","违反平台","严重违规","请勿发布","不当内容","无法协助生成","平台安全"
Table 7: A list of keywords used to identify and filter out model refusals or generic non-answers during data preprocessing.
Synthetic Data Generation Template
违规类别选项 = ["博彩", "低俗色情", "谩骂引战", "欺诈", "黑产广告","不违规"]
性别选项 = ["男性", "女性", "未知"]
年龄选项 = ["0-12", "13-17", "18-24", "25-34", "35-44", "45-54", "55-64", "65-74", "75-84", "85+"]
职业选项 = [ "程序员", "软件工程师", "系统分析师", "网络安全工程师", "数据分析师", "人工智能工程师", "硬件工程师", "数据库管理员", "前端开发", "后端开发", "教师", "大学教授", "辅导员", "培训师", "教研员", "教务管理", "保育员", "早教指导师", "在线教育运营", "教育产品经理", "医生", "护士", "药剂师", "营养师", "心理咨询师", "康复治疗师", "兽医", "急救员", "公共卫生管理", "医学研究员", "律师", "法官", "检察官", "公证员", "法律顾问", "公务员", "警察", "消防员", "海关人员", "边检人员", "工程师", "机械工程师", "电气工程师", "土木工程师", "化工工程师", "建筑设计师", "施工监理", "测绘员", "工业机器人操作", "质量检测员", "销售", "市场", "市场营销", "品牌策划", "广告策划", "产品经理", "客户经理", "渠道经理", "采购专员", "供应链管理", "设计师", "平面设计师", "室内设计师", "服装设计师", "插画师", "摄影师", "影视编导", "配音演员", "游戏原画师", "舞台美术", "厨师", "酒店经理", "导游", "空乘人员", "健身教练", "美容师", "美发师", "按摩技师", "客服专员", "速递员", "粮农", "菜农", "果农", "猪农", "渔民", "牧民", "林业员", "园艺师", "农业技术员", "水产养殖", "自媒体运营", "直播主播", "电竞选手", "无人机飞手", "碳排放管理员", "陪诊师", "收纳师", "研学旅行指导", "宠物殡葬师", "民宿管家", "工人", "个体经营者", "保安", "司机", "维修工", "电工", "木工", "搬运工", "环卫工", "门卫", "未知" ]
学历选项 = ["文盲", "小学", "初中", "高中", "中专", "大专", "本科", "硕士", "博士", "未知"]
长度范围选项 = ["0-5", "6-10", "11-15", "16-20", "21-25", "26-30", "31-35", "36-40", "41-45", "46-50", "50+"]
平台选项 = ["微博", "小红书", "QQ", "微信", "抖音", "B站", "知乎", "快手", "豆瓣", "百度贴吧"]
规避手段选项 = ["拼音", "谐音词", "形似词", "emoji", "不规避"]
规避策略说明选项 = [ "该文本使用关键词替换策略,使用{规避手段}来规避部分敏感词汇或内容。", "该文本为正常文本,没有使用任何规避策略或手段。" ]
叙事角度选项 = ["第一人称", "第二人称", "第三人称"]
【角色设定】 你是一位经验丰富的互联网内容运营专家,擅长模仿不同群体的真实语言特征。请根据以下参数生成文本: 【用户画像】 - 性别:{性别} - 年龄:{年龄} - 职业:{职业} - 学历:{学历} 【文本建模】 - 是否违规:{是否违规} - 违规类别:{违规类别} - 违反规则:{规则} - 文本长度:{文本长度}字 - 叙事角度:{叙事角度} - 发布平台:{平台} 【规避策略】 - 规避手段:{规避手段} - 说明:{规避策略说明} 【生成要求】 请生成符合用户画像和文本建模要求的内容。如果使用了规避策略,请注意正确应用,以确保内容尽可能真实。请直接输出生成的文本。
Table 8: Structured prompt template and attribute options for synthetic illicit content generation. The template covers violation categories, persona features, text features, evasion tactics, and strategy explanations, enabling comprehensive simulation of real-world harmful content detection scenarios.